# Temporal Knowledge-Graph Memory in a Partially Observable Environment

### v3

@@ 3 Room Environment v3
(a) Schematic
(b) Knowledge-graph view
Figure 1: Two views of the same hidden state ($s_{t=99}$). The agent’s observation ($o_{t}$) is the induced RDF subgraph of the current room and its visible adjacency relations. Full step-by-step visualizations and convergence analysis are provided in the anonymous repository.
The Room Environment v3 (Figure 1) is a configurable $g\!\times\!g$ grid world populated with named rooms, static and moving objects, and periodically changing inner walls. All environment components—hidden dynamics, observations, and the QA loop—are defined in RDF triples triples, making the environment a minimal but fully KG-shaped testbed for studying temporal memory.
@@ 7 Conclusion
8 Acknowledgments
This research was (partially) funded by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organization for Scientific Research, https://www.hybrid-intelligence-centre.nl/.

### v4

@@ 3 Room Environment v3
Figure 1: Bird’s-eye schematic of the hidden state at $t=99$ ($s_{t=99}$), showing spatial layout and entity placement. This view is only a schematic for intuition: the actual environment state and agent-facing world are represented in the RDF knowledge-graph world (Figure 2). The agent does not directly observe $s_{t}$; its observation $o_{t}$ is the induced RDF subgraph of the current room and visible adjacency relations.
Figure 2: Knowledge-graph view of the same hidden state at $t=99$ ($s_{t=99}$), expressed as RDF structure.
The Room Environment v3 (Figures 1 and 2) is a configurable $g\!\times\!g$ grid world populated with named rooms, static and moving objects, and periodically changing inner walls. All environment components—hidden dynamics, observations, and the QA loop—are defined in RDF triples triples, making the environment a minimal but fully KG-shaped testbed for studying temporal memory.